OpenStudy (anonymous):

A theater has 20 rows of chairs. The first row has 16 chairs, the second has 18, the third has 20, and so on. In total, the theater seats _____ people.

OpenStudy (anonymous):

a=16 d=18-16=2 n=20 \[S _{_{n}}=\frac{ n }{ 2 }\left\{ 2a+\left( n-1 \right)d \right\}\] ?